OpenStudy (anonymous):

Write a formula for this sequence: 8, -1, -10, -19,...

OpenStudy (anonymous):

It degrees by 9 each time, I call this the slope m = -9 If there was a number before the 8, it would be 8 + 9 = 17, I call this the y-intercept so we have y = -9x + 17 where x represents the placement of the term and y represents the term value

OpenStudy (anonymous):

y=-9x+17, same here xD

OpenStudy (anonymous):

any easy logic?

OpenStudy (anonymous):

did you see my slope logic?

OpenStudy (anonymous):

ya,...but it is graphical

OpenStudy (anonymous):

Sequences: Formula: Un=U1+(n-1)d U1=8 d=-9 Un=8-9(n-1) Un=-9n+17 If you don't understand the formula, maybe try looking at this: http://jwilson.coe.uga.edu/EMAT6680Fa06/Crumley/Sequences/Sequences1.html

OpenStudy (anonymous):

hey thats easy

OpenStudy (anonymous):

was the explanation not needed? :S sry if it was :(

OpenStudy (anonymous):

hey no.....u diid help a lot

OpenStudy (anonymous):

oh, ok :) that's good